### v1

# A phase-space view of vibrational energies without the Born-Oppenheimer framework

Abstract
We show that following the standard mantra of quantum chemistry and diagonalizing the Born-Oppenheimer (BO) Hamiltonian $\hat{H}_{\rm BO}({\bm{R}})$ is not the optimal means to construct potential energy surfaces. A better approach is to diagonalize a phase-space electronic Hamiltonian, $\hat{H}_{\rm PS}({\bm{R}},{\bm{P}})$, which is parameterized by both nuclear position ${\bm{R}}$ and nuclear momentum ${\bm{P}}$. The foundation of such a non-perturbative phase-space electronic Hamiltonian can be made rigorous using a partial Wigner transform and the method has exactly the same cost as BO for a semiclassical calculation (and only a slight increase in cost for a quantum nuclear calculation). For a three-particle system, with two heavy particles and one light particle , numerical results show that a phase-space electronic Hamiltonian produces not only meaningful electronic momenta (which are completely ignored by BO theory) but also far better vibrational energies. As such, for high level results and/or systems with degeneracies and spin degrees of freedom, we anticipate that future electronic structure and quantum chemistry packages will need to take as input not just the positions of the nuclei but also their momenta.
I Introduction
I.1 The Hydrogen Atom
The Born-Oppenheimer (BO) framework[1, 2, 3] is the standard approach to understanding chemistry, upon which the vast majority of chemical intuition is built. The basic ansatz of BO theory is that, because nuclei are slow and electrons are fast, one can separate the total Hamiltonian into two terms, the nuclear kinetic energy and the electronic Hamiltonian:
$\hat{H}=\frac{{\hat{\bm{P}}}^{2}}{2M}+\hat{H}_{\rm el}(\hat{\bm{R}}).$ (1)
One diagonalizes $\hat{H}_{\rm el}(\hat{\bm{R}})$ in order to construct potential energy surfaces, and then the BO approximation stipulates that nuclear motion follows a single surface.
One critical limitation of BO theory can be identified by considering the hydrogen atom, the most well understood problem in quantum mechanics. The Hamiltonian is simply the kinetic energy of the electron plus the kinetic energy of the proton plus their interaction:
$\hat{H}_{\rm Hydrogen}=\frac{\hat{\bm{P}}_{\rm n}^{2}}{2M}+\frac{\hat{\bm{p}}_{\rm e}^{2}}{2m}-\frac{1}{|{\bm{\hat{r}}_{\rm e}-\bm{\hat{R}}_{\rm n}}|}.$ (2)
As can be found in any textbook[4], the way to solve this Hamiltonian is to change variables from $\hat{\bm{R}}_{\rm n}$ and $\hat{\bm{r}}_{\rm e}$ to the relative position and center of mass coordinates:
$\hat{\bm{r}}=\hat{\bm{r}}_{\rm e}-\hat{\bm{R}}_{\rm n},\quad\hat{\bm{R}}=\frac{M\hat{\bm{R}}+m_{\rm e}\hat{\bm{r}}}{M+m_{\rm e}}.$ (3)
In so doing, the Hamiltonian becomes:
$\hat{H}_{\rm com}=\frac{\hat{\bm{P}}^{2}}{2(m_{\rm e}+M)}+\frac{\hat{\bm{p}}^{2}}{2\mu}-\frac{1}{|\hat{\bm{r}}|}$ (4)
with the reduced mass $\mu=(M^{-1}+m_{e}^{-1})^{-1}$, and the stationary energies for the system are the center of mass kinetic energy plus the quantized bound state energy of the atom:
$\epsilon_{n}^{\rm exact}(\bm{P})=\frac{\bm{P}^{2}}{2(M+m_{\rm e})}-\frac{\mu e^{4}}{8\epsilon_{0}^{2}h^{2}}\frac{1}{n^{2}}.$ (5)
Note that, within the BO (clamped nuclei) approximation, the result is slightly different if one diagonalizes $\hat{H}_{el}$:
$\epsilon_{n}^{\rm BO}(\bm{P})=\frac{\bm{P}^{2}}{2M}-\frac{m_{\rm e}e^{4}}{8\epsilon_{0}^{2}h^{2}}\frac{1}{n^{2}}.$ (6)
Now, on the one hand, one might argue that the energy expressions are not that different between Eq. 5 and Eq. 6. That being said, if we ignore energy for a moment, there is a conceptual problem insofar as the fact that, unless one changes coordinates, the electronic momentum of the ground state must be zero[5, 6, 7]; in other words, even if the hydrogen atoms is moving, the electronic momentum is still calculated to be zero (which is incorrect). Moreover, in the context of energy, if we were to consider a system of $N$ isolated hydrogen atoms, the total difference in energy becomes $N$ times larger. Furthermore, if one were to bring those $N$ atoms closer together, so that they could interact and form hydrogen molecules or solids, there is no reason to assume that the problem of freezing nuclei (rather than choosing optimal coordinates) would go away[8, 9, 10, 11, 12, 13, 14] – in fact, one might expect that with more and more interacting nuclei, the problems of BO theory would only become more daunting and more difficult to solve.
Within the context of molecular quantum chemistry, the problem above is often addressed in the literature by separating out the center of mass, in the spirit of Eq. 4 above[15, 16, 17]. Mathematically, one can formally change coordinates and separate out the total center of mass by choosing Jacobi coordinates for the nuclei and relating all electrons to the nuclear center of mass:
$\hat{\bm{R}}_{\rm n,com}=\frac{\sum_{A}M_{A}{\hat{\bm{R}}_{A}}}{\sum_{A}M_{A}},\quad\hat{\bm{r}}_{i}=\hat{\bm{r}}_{i}-\hat{\bm{R}}_{\rm n,com}.$ (7)
It is crucial to emphasize that these transformations do not solve the problem above – especially in the context of condensed phase chemistry with many molecules. In particular, even for a pair of H atoms separated far from each other, e.g. one on the earth and one on the moon, the transformation in Eq. 7 introduces one coordinate (the nuclear center of mass) which is somewhere in the exosphere, but does not succeed at recovering the correct ground state energy.
I.2 Review of Electronic Phase-Space Hamiltonians
New approaches that go beyond BO theory are clearly needed for accurate quantum chemistry[18, 19, 20]. With this in mind, we have recently proposed a phase-space approach to treat the exact problem above. To motivate this approach, consider the Hamiltonian in Eq. 2 as rewritten in an atomic adiabatic basis (i.e. after diagonalization of $\hat{U}\hat{H}_{\rm el}\hat{U}^{\dagger}=\hat{V}_{\rm ad}$):
$\hat{H}_{\rm ad}=\frac{\hat{\bm{P}}_{\rm n}^{2}}{2M}+\hat{V}_{\rm ad}-\frac{\hat{\bm{P}}_{\rm n}\cdot\hat{\bm{p}}_{\rm e}^{\rm ad}+\hat{\bm{p}}_{\rm e}^{\rm ad}\cdot\hat{\bm{P}}_{\rm n}}{2M}+\frac{(\hat{\bm{p}}_{\rm e}^{\rm ad})^{2}}{2M}.$ (8)
where $\hat{\bm{p}}_{\rm e}^{\rm ad}=\hat{U}\hat{\bm{p}}_{e}\hat{U}^{\dagger}$. If one now makes the ansatz that the nuclei are classical, one can construct a phase-space electronic Hamiltonian which is a function of the parameters (not operators) $\bm{R}_{\rm n}$ and $\bm{P}_{\rm n}$ and transform it back to the original basis:
$\hat{H}_{\rm PS}=\frac{{\bm{P}}_{\rm n}^{2}}{2M}+\hat{H}_{\rm el}-\frac{{\bm{P}}_{\rm n}\cdot\hat{\bm{p}}_{\rm e}}{M}+\frac{\hat{\bm{p}}_{\rm e}^{2}}{2M}.$ (9)
A simple analysis shows that $\hat{H}_{\rm PS}$ represents a boost of the electron with nuclear momentum ${\bm{P}}_{\rm n}$, and its corresponding eigenvalues are identical to the exact energies for a hydrogen atom:
$\epsilon_{n}^{\rm PS}({\bm{P}}_{\rm n})=\frac{\bm{P}_{\rm n}^{2}}{2M}-\frac{\mu e^{4}}{8\epsilon_{0}^{2}h^{2}}\frac{1}{n^{2}}$ (10)
(provided we equate ${\bm{P}}_{\rm n}$ [which captures the nuclear momentum] with $\bm{P}$ [which is formally the center of mass momentum]).
This finding is very encouraging because it offers an approach for a system with $N$ isolated hydrogen atoms. In such a case, all we must do is to partition the electronic momentum into different operators, one around each atom, so that one can effectively perform $N$ different local changes of coordinate (around each atom):
$\hat{\bm{p}}\rightarrow\frac{1}{2}\sum_{A}\left(\hat{\theta}_{A}\hat{\bm{p}}+\hat{\bm{p}}\hat{\theta}_{A}\right)$ (11)
where $\hat{\theta}_{A}(\hat{\bm{r}})$ is a partition of space:
$\sum_{A}\hat{\theta}_{A}(\hat{\bm{r}})=1.$ (12)
Note that while there are many ways to choose the space partition function $\hat{\theta}_{A}$, a natural option is to choose smooth functions that are localized on atoms (e.g., a combination of Gaussian functions, see supporting information [SI]).
One can then write:
$\hat{H}_{\rm PS}$ $=$ $\sum_{A}\frac{{\bm{P}_{A}}^{2}}{2M_{A}}-\frac{1}{2}\sum_{A}\frac{\bm{P}_{A}}{M_{A}}\cdot\left(\hat{\theta}_{A}\hat{\bm{p}}+\hat{\bm{p}}\hat{\theta}_{A}\right)$ (13)
$+$ $\frac{1}{4}\sum_{A}\frac{\left(\hat{\theta}_{A}\hat{\bm{p}}+\hat{\bm{p}}\hat{\theta}_{A}\right)^{2}}{2M_{A}}+\hat{H}_{\rm el}({\bm{R}})$
or equivalently
$\hat{H}_{\rm PS}$ $=$ $\sum_{A}\frac{{\left(\bm{P}_{A}-i\hbar\hat{\bm{\Gamma}}_{A}^{\prime}\right)}^{2}}{2M_{A}}+\hat{H}_{\rm el}({\bm{R}})$ (14)
with
$\hat{\bm{\Gamma}}_{A}^{\prime}=\frac{1}{2i\hbar}\left(\hat{\theta}_{A}\hat{\bm{p}}+\hat{\bm{p}}\hat{\theta}_{A}\right).$ (15)
This is almost the final form of $\hat{H}_{\rm PS}$, but there are two more important points. First, one can show that if one diagonalizes Eq. 14, then classical dynamics propagated along the resulting energy surfaces will automatically conserve linear momentum. If one wishes to conserve angular momentum, one must also include another term to account for angular momentum[21]:
$\hat{H}_{\rm PS}$ $=$ $\sum_{A}\frac{{\left(\bm{P}_{A}-i\hbar\hat{\bm{\Gamma}}_{A}^{\prime}-i\hbar\hat{\bm{\Gamma}}_{A}^{\prime\prime}\right)}^{2}}{2M_{A}}+\hat{H}_{\rm el}({\bm{R}})$ (16)
A preliminary analysis of the relevant equations of motion is given in the SI.
Second, the form of the phase-space Hamiltonian in Eq. 14 is reasonable insofar as, if one diagonalizes $\hat{H}_{\rm el}(\bm{R})$, the exact Hamiltonian in the basis of electronic eigenstates is of the form:
$\hat{H}_{\rm ad}=\sum_{A}\frac{{\left(\hat{\bm{P}}_{A}-i\hbar\hat{\bm{d}}_{A}\right)}^{2}}{2M_{A}}+\hat{V}_{\rm ad}(\hat{\bm{R}})$ (17)
where $\hat{\bm{d}}_{A}$ is the derivative coupling and $\hat{V}_{\rm ad}$ is the diagonal matrix of eigenvalues. Thus, one concludes that the phase-space Hamiltonian is the natural electronic object that emerges when we make the nuclear position and momentum classical,
$\hat{\bm{R}}\rightarrow\bm{R},\quad\hat{\bm{P}}\rightarrow\bm{P}$ (18)
and one further approximates the derivative coupling by an operator $\hat{\bm{d}}_{A}\rightarrow\hat{\bm{\Gamma}}_{A}=\hat{\bm{\Gamma}}^{\prime}_{A}+\hat{\bm{\Gamma}}^{\prime\prime}_{A}$. By construction, $\hat{\bm{\Gamma}}_{A}$ satisfies[21, 22, 23]
$-i\hbar\sum_{A}\hat{\bm{\Gamma}}_{A}+\hat{\bm{p}}$ $=$ $0,$ (19)
$\Big{[}-i\hbar\sum_{B}\frac{\partial{}}{\partial{\bm{R}_{B}}}+\hat{\bm{p}},\hat{\bm{\Gamma}}_{A}\Big{]}$ $=$ $0,$ (20)
$-i\hbar\sum_{A}{\bm{R}}_{A}\times\hat{\bm{\Gamma}}_{A}+\hat{\bm{l}}+\hat{\bm{s}}$ $=$ $0,$ (21)
$\Big{[}-i\hbar\sum_{B}\left(\bm{R}_{B}\times\frac{\partial{}}{\partial{\bm{R}_{B}}}\right)_{\gamma}+\hat{l}_{\gamma}+\hat{s}_{\gamma},\hat{\Gamma}_{A\delta}\Big{]}$
$=i\hbar\sum_{\alpha}\epsilon_{\alpha\gamma\delta}\hat{\Gamma}_{A\alpha},$ (22)
and thus we may conclude that the $\hat{\bm{\Gamma}}^{\prime}$ and $\hat{\bm{\Gamma}}^{\prime\prime}$ operators are akin to the local translational and rotational components of the derivative couplings[24, 25, 26, 27, 23], so that the phase-space Hamiltonian in Eq. 14 is an approximate (but tractable) version of Eq. 17.
I.3 Our Goal In this Article
This concludes our heuristic review of phase-space electronic Hamiltonian theory. We have sought to justify the phase-space approach by appealing to the readers’ understanding of the importance of a change of coordinates when simulating a simple hydrogen atom, but have also demonstrated that the approach is reasonable by mapping the approximate $\hat{\bm{\Gamma}}$ factors to the derivative coupling. By this point in time, there is already ample evidence in the literature that this phase-space approach can recover electronic momenta[22], electronic current densities[23] and even vibrational circular dichroism (VCD) signals[28]. In all cases, we have found that a phase-space electronic Hamiltonian is simply more accurate, and offers more and richer set of information than a BO Hamiltonian.
That being said, however, one could argue that we still do not have a smoking gun that clearly establishes the supremacy of the phase-space approach. Furthermore, one might also note that the phase-space approach developed in Sec. I.2 does require some hand-waving, and for complete confidence, one would like to understand how to derive and work with such a Hamiltonian in the context of an exact calculation. These concerns motivate the present article. In what follows, we will first establish how a phase-space electronic Hamiltonian can be rigorously defined through a Wigner transformation, and then, using the inverse Weyl transformation, we will show how to extract quantized vibrational energies from a phase-space approach. In particular, we will show that vibrational energies extracted by diagonalizing the nuclear Hamiltonian from a phase-space electronic Hamiltonian universally outperforms (i.e. gives more accurate vibrational energies) than diagonalizing the corresponding nuclear Hamiltonian from a Born-Oppenheimer electronic Hamiltonian. This finding would appear to be the smoking gun that makes clear that, in the future, quantum chemistry packages will need to be prepared to avoid the Born-Oppenheimer Hamiltonian.
II Born-Oppenheimer versus phase-space theory of vibrational energies
Before introducing phase-space theory, it will be helpful to review briefly the easiest way to derive BO theory from a formal point of view.
II.1 Standard BO theory
The usual BO framework proceeds by initially defining a unitary operator $\hat{U}$ that diagonalizes $\hat{H}_{\rm el}$. One then transforms to the new adiabatic basis,
$\hat{H}_{\rm ad}$ $=$ $\hat{U}\hat{H}\hat{U}^{\dagger}=\frac{(\hat{\bm{P}}-i\hbar\hat{\bm{d}})^{2}}{2M}+\hat{V}_{\rm ad}(\hat{\bm{R}}).$ (23)
The BO assumption is then that one can neglect the $\hat{\bm{P}}\cdot\hat{\bm{d}}$ and $\hat{\bm{d}}\cdot\hat{\bm{d}}$ terms in Eq. 17, and diagonalize:
$\hat{H}_{\rm BO}=\frac{\hat{\bm{P}}^{2}}{2M}+\hat{V}_{\rm ad}(\hat{\bm{R}}).$ (24)
Note that, because $\hat{V}_{\rm ad}={\rm diag},(\epsilon_{1}(\hat{\bm{R}}),\cdots,\epsilon_{n}(\hat{\bm{R}}))$ is diagonal, this amounts to many different diagonalizations of $\frac{\hat{\bm{P}}^{2}}{2M}+\epsilon_{n}({\bm{R}})$. Note also that, while $\hat{H}_{\rm ad}$ is a unitary transform of $\hat{H}$, $\hat{H}_{\rm BO}$ is clearly not such a transform.
Everything above is well known and standard. Also well known, but slightly less standard, is that the BO approximation can also be derived through partial Wigner transforms[29, 30, 31]. As a reminder, the partial Wigner transform ($\mathcal{W}$) for any operator $\hat{O}$ is given by:
$\hat{O}_{W}({\bm{R}},{\bm{P}})$ $=$ $\mathcal{W}\left[\hat{O}\right]$ (25)
[…]
where $\hat{O}_{W}({\bm{R}},{\bm{P}})$ is an matrix of electronic operators that depends parametrically on ${\bm{R}}$ and ${\bm{P}}$. This transformation is invertible by a Weyl transform, and one can show that:
$\hat{O}$ $=$ $\mathcal{W}^{-1}\left[\hat{O}_{W}\right],$ (27)
with matrix elements
$\bra{\bm{R}}\hat{O}\ket{\bm{R}^{\prime}}$ $=$ $\int\frac{d\bm{P}}{2\pi\hbar}e^{\frac{i}{\hbar}\bm{P}\cdot(\bm{R}-\bm{R}^{\prime})}\hat{O}_{W}\left(\frac{\bm{R}+\bm{R}^{\prime}}{2},P\right).$ (28)
Now, to formally derive the BO approximation, we build and diagonalize the Wigner transform of $\hat{H}$:
$\hat{H}_{W}$ $=$ $\frac{\bm{P}^{2}}{2M}+\hat{H}_{\rm el}(\bm{R})=\hat{U}_{W}^{\dagger}\hat{\Lambda}_{W}\hat{U}_{W}.$ (29)
The BO Hamiltonian in Eq. 24 is then exactly equivalent to the Weyl transform of $\Lambda_{W}$:
$\hat{H}_{\rm BO}=\mathcal{W}^{-1}\left[\hat{\Lambda}_{W}\right]=\mathcal{W}^{-1}\left[\hat{U}_{W}\hat{H}_{W}\hat{U}_{W}^{\dagger}\right].$ (30)
Note that the BO Hamiltonian is (obviously) not a unitary transform of the exact Hamiltonian. If one were to apply a Wigner transformation to the adiabatic Hamiltonian in Eq. 23, one must remember that the Wigner transform of a product of operators is not the product of Wigner transforms but rather involves a Moyal product between them. In such a case, it follows that the exact Hamiltonian can be written as:
$\hat{H}_{\rm ad}$ $=$ $\mathcal{W}^{-1}\left[\hat{U}_{W}*\hat{H}_{W}*\hat{U}_{W}^{\dagger}\right]=\mathcal{W}^{-1}\left[\left(\hat{U}\hat{H}\hat{U}^{\dagger}\right)_{W}\right].$ (31)
Observe that the difference between Eq. 30 and Eq. 31 is that normal product multiplication is replaced by Moyal star products.
In any event, it is quite common today to extract vibrational energies from a BO Hamiltonian. To do so, for the ground state, one simply diagonalizes the nuclear Hamiltonian for the electronic ground state $\bra{\psi_{0}}\hat{H}_{\rm BO}\ket{\psi_{0}}$ (which is equivalent to diagonalizing $\frac{\hat{\bm{P}}^{2}}{2M}+\epsilon_{0}(\hat{\bm{R}})$). Alternatively, within a harmonic approximation, one can can simply expand $\epsilon_{0}$ (or equivalently $\bra{\psi_{0}}\hat{\Lambda}_{W}\ket{\psi_{0}}$) to second order in $\bm{R}$ and apply the standard harmonic oscillator result[32].
II.2 Phase-Space Electronic Hamiltonian
According the theory presented in Sec. I.2, there are better means to approximate the full Hamiltonian than $\hat{H}_{\rm BO}$. In generating such approximations, and in light of Eq. 16, it would appear that the optimal approach is to parameterize the approximate Hamiltonian by both nuclear position and momentum. Rigorously, such a parameterization can only occur after a Wigner transformation. Thus, the phase-space approach is to construct an electronic Hamiltonian of the form:
$\hat{H}_{W}^{\rm PS}$ $=$ $\hat{H}_{W}-i\hbar\hat{\bm{\Gamma}}_{W}\cdot\frac{\bm{P}}{M}-\hbar^{2}\frac{\hat{\bm{\Gamma}}_{W}^{2}}{2M},$ (32)
and then diagonalize it:
$\hat{H}^{\rm PS}_{W}$ $=$ $L_{W}^{\dagger}\hat{\Lambda}^{\rm PS}_{W}L_{W}.$ (33)
The final phase-space electronic Hamiltonian is then constructed (as for BO) by a Weyl transformation:
$\hat{H}_{\rm PS}$ $=$ $\mathcal{W}^{-1}\left[\Lambda_{W}^{\rm PS}\right]=\mathcal{W}^{-1}\left[L_{W}H_{W}^{\rm PS}L_{W}^{\dagger}\right]$ (34)
Eq. 34 should be compared with Eq. 30. Note that, just like $\hat{H}_{\rm BO}$, $\hat{H}_{\rm PS}$ is not a unitary transform of $\hat{H}$. Also, just as for $H_{\rm BO}$, vibrational energies can be formally extracted by diagonalizing the nuclear Hamiltonian $\bra{\psi_{0}^{\rm PS}}\hat{H}_{\rm PS}\ket{\psi_{0}^{\rm PS}}$, but in this case, constructing $\hat{H}_{\rm PS}$ requires a Weyl transform in Eq. 34. Put differently, while working with the phase-space electronic Hamiltonian is obvious for semiclassical simulations, for quantum nuclear (i.e. vibrational) problems, where ${\bm{P}}$ is an operator (not a paramter), the phase-space electronic Hamiltonian has no simple analogue to Eq. 24 (but rather requires a Weyl transformation).
As discussed above, previous work has demonstrated that a phase-space approach has many advantages over a BO Hamiltonian – one recovers the exact eigenvalues of a hydrogen atom, one recovers nonzero electronic momenta and electronic current density, and one can also match experimental VCD spectroscopy. In what follows, we will now show that, using the Hamiltonian in Eq. 34, we also recover improved vibrational energies (often much improved).
III Demonstration through a model problem
To demonstrate the power of a phase-space approach for calculating vibrational spectra, consider a classic model Hamiltonian that describes a one-dimensional asymmetric hydrogen bond system $\rm O-H\cdots O$ from Refs.33, 14. In this coarse-grained model, the light particle represents a hydrogen atom with mass $m$, and the heavy particles are oxygen atoms with mass $M$ – though we will also refer to the particle as an electron and the heavy particle as a nucleus, given the importance of BO theory to electronic structure calculations. The potential from Ref.33, 14 is plotted in Fig. 1(a) as a function of the heavy and light degrees of freedom (DoF) $R$ and $r$, where $R$ is the distance between the two oxygen atoms and the light DoF $r$ indicates the distance between the H atom to the center of two O atoms. The intuition behind this model Hamiltonian is that the hydrogen will stay at the center when the two oxygen atoms are close, and it will bonded to one oxygen when the oxygen atoms are far away due to the hydrogen bonding effects. All of the relevant details of the Hamiltonian can be found in SI.
For this model Hamiltonian, we have calculated vibrational data using three different approaches:
  1. 1.
    Exact calculations, where we diagonalize $\hat{H}$ over a joint grid in the nuclear coordinate $R$ and the electronic coordinate $r$.
  2. 2.
    BO calculations, where we loop over a grid of nuclear positions $R_{i}$, use a grid of electronic positions to diagonalize the electronic Hamiltonian associated to each nuclear grid point $\hat{H}_{\rm el}(R_{i})\ket{\psi_{0}}=\epsilon_{0}(R_{i})\ket{\psi_{0}}$, and then diagonalize the nuclear Hamiltonian $\hat{P}^{2}/2M+\epsilon_{0}(\hat{R})$ over the nuclear grid.
  3. 3.
    PS calculations, where we loop over a joint grid of nuclear positions and momenta $(R_{i},P_{j})$, use a grid of electronic positions to diagonalize the phase-space Hamiltonian (Eq. 32) associated to each joint grid point $\hat{H}_{W}^{\rm PS}(R_{i},P_{j})\ket{\psi_{0}}=\epsilon^{PS}_{W,0}(R_{i},P_{j})\ket{\psi_{0}}$, invert $\epsilon^{PS}_{W,0}(R_{i},P_{j})$ to a function in real space $\epsilon_{0}^{\rm PS}(R_{i},R_{j})$ using Eq. 28, and then diagonalize $\epsilon_{0}^{PS}(R_{i},R_{j})$ over the nuclear grid.
Below, we will also report results where we consider harmonic BO and PS surfaces, i.e. where we force $\epsilon_{0}(R)$ or $\epsilon^{\rm PS}_{W,0}(R,P)$ to be a quadratic function. More numerical details for these simulations can be found in SI.
To begin our analysis, in Fig. 1(b), we plot the ground BO potential energy surface that arises from treating the light and heavy degrees of freedom (DoF) separately, i.e., treating the oxygen atoms as nuclei and the hydrogen atom as an electron. One should not expect this BO separation will not be a good approximation if the mass ratio $M/m$ is not large enough, such as the actual masses of oxygen and hydrogen atoms (16 and 1 atomic mass units, respectively). Thus, by varying the nuclei/electron mass ratio $M/m$, we can control non-adiabatic effects for this model Hamiltonian.
Figure 1: (a) Two-dimensional potential energy surface as a function of inter-nuclear distance ($R$) and distance between electron to the nuclear center of mass ($r$) (b) Morse-like BO potential energy surface obtained from diagonalizing the electronic Hamiltonian as a function of the nuclear position $R$.
In Fig. 2, we report our results for the ground to first excited vibrational energy gap as a function of the different mass ratio. For all methods, the absolute values of energy gap decreases as the mass ratio increases and converge to the same results in the large mass ratio limit ($\sim$1000 corresponding to a typical nuclear/electronic mass ratio) in Fig. 2(a). Thus, in Fig. 2(b), we plot the error of the energy gap as compared to the exact results. In the small mass ratio limit ($\sim$10 corresponding to the oxygen and hydrogen atoms), one can clearly identify that the phase-space method shows a much smaller error compared to the BO method. Notably, even within a harmonic approximation, the phase-space method still outperforms the BO and sometimes can be comparable to the quantum BO results. Lastly, to best understand these results, the inset of Fig. 2(b) shows the relative errors in the logarithm scale. As one might expect, the phase-space and BO energy errors both scale with the same power law ($\sim m/M$), but the phase-space method consistently gives an order of magnitude smaller relative error even in the large mass ration limit.
Figure 2: (a) The ground to first excited vibrational energy gap as calculated using various methods. (b) The error in the vibrational energy gap as compared to the exact results. Phase-space methods have far smaller error as compared to BO methods across the entire mass ratio range. (Inset: The relative error on a logarithmic scale.)
It must be emphasized that, with a phase-space Hamiltonian, one can extract other useful properties as well (besides energy) and these quantities are also more accurate than the corresponding BO Hamiltonian. For example, in Fig. 3, we show the errors of expectation values of a few nuclear and electronic observables. First, for the electronic momentum, it is important to note that BO calculations always predict zero, as $\bra{\psi_{k}}\hat{\bm{p}}_{\rm e}\ket{\psi_{k}}=0$ for all BO electronic states at any nuclear geometry. However, as shown in Fig. 3(a), the phase-space method can capture the electronic momentum between the ground and first excited vibrational states with quantitative accuracy. Second, for nuclear momentum, Fig. 3(b) demonstrates that phase-space results are also much more accurate than BO methods. One interesting nuance is that one must replace the nuclear momentum operator $\hat{\bm{P}}$ by $\hat{\bm{P}}-\hat{\bm{\Gamma}}$ in the phase-space formalism which represents a canonical – kinetic momentum correspondence. Overall, for the off-diagonal electronic and nuclear momentum (between the same nuclear vibrational eigenstates), the phase-space method performs very well even with the harmonic approximation. As a side note, the on-diagonal nuclear and electronic momentum always vanish due to time reversal symmetry of the whole system (nuclei + electron). Finally, in Fig. 3(c) and (d), while the nuclear and electronic position observables predicted by BO and phase-space methods are not very different, the phase-space methods do provide slightly better results for the excited vibrational states. For further analysis, and a decomposition of the importance of different terms in the phase-space Hamiltonian, see the SI.
Figure 3: The errors of various observables as calculated from BO and phase-space methods and compared against exact results. We plot the errors in (a) the electronic momentum and (b) the nuclear momentum between the ground and first excited nuclear vibrational state, where the phase-space approach are strong improvements to the BO results. We also plot the errors in (c) electronic position and (d) nuclear position for the ground and first excited vibrational states, where the phase-space and BO results are similar.
IV Discussion and Conclusions
In this work, we have demonstrated that a novel phase-space view of electronic structure theory can vastly outperform Born-Oppenheimer theory in many facets. Not only can the phase-space approach deliver nonzero and accurate electronic momentum matrix elements between ground state vibrational states (which is impossible for Born-Oppenheimer theory), but the method can also predict molecular vibrational energies that are far more accurate than those produced by BO theory – especially when the mass difference between the light and heavy particles is not too large.
Looking forward, it would appear clear from the plethora of data collected here and in Refs.22, 23 that a phase-space approach is more accurate than a BO approach – and with roughly the same cost. That being said, one must wonder: in what circumstances will the phase-space approach yield meaningfully different answers than BO theory? To that end, one can imagine several exciting future problems where one would like to go beyond the limitation of BO theory. The first potential arena would be to study proton transfer[34] and proton coupled electron/energy transfer[35, 36] where one could combine a phase-space approach with electronic structure methods that treat both light nuclei (protons) and electrons quantum mechanically (e.g., multi-component methods[37, 38, 39] and nuclear-electronic orbital methods (NEO)[40, 41]. Second, it will be fruitful to apply a phase-space approach to study magnetic molecules and solids, where spin-symmetry breaking can lead to intrinsic chiral phonon modes and alter vibrational frequencies[42, 43], and electronic momentum is known to be critically important in such circumstances[44, 45]. Third, given the ab initio and semiclassical nature of the phase-space electronic Hamiltonian, one immediately goal is to investigate nuclear (vibrational) effects at curve crossings during electron transfer processes[46] and decipher whether phase-space approaches (that include angular momentum conservation[47]) can help explain the nature of nuclear effects in chiral-induced spin selectivity[48, 49, 50]. Fourth, the present phase-space approach opens new and inexpensive avenues for exploring electron-phonon interactions, which could lead to a better atomistic understanding of superconductivity[51, 52, 53]. At the end of the day, there is a very real chance that phase-space electronic Hamiltonian approaches are here to stay and that a new world of very novel chemistry awaits us.